OpenStudy (anonymous):

Find percent notation (Simplify answer) \[\frac{ 17 }{ 25 }\]

jimthompson5910 (jim_thompson5910):

hint: multiply top and bottom by 4 we multiply by 4 since 4*25 = 100

OpenStudy (anonymous):

\[\frac{ 68 }{ 100 }\]

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

so it would be 68% ?

jimthompson5910 (jim_thompson5910):

anything over 100 is equal to that percentage therefore, 68/100 = 68%

jimthompson5910 (jim_thompson5910):

yep you got it

OpenStudy (anonymous):

alright thanks

jimthompson5910 (jim_thompson5910):

you're welcome